in direct variation 2y=3x, what is the K value?
a)3/2
b)3
c)2/3
d)2
what is the slope of the direct variation y=3/5x
a)3
b)5
c)3/5
d)5/3

Answer by Mathtut(3670)

does x vary directly with y or vice versa
:
if x varies directly as y then 3x=2y--->x/y=2/3.....so k=2/3
if y varies directly as x then 2y=3x--->y/x=3/2.....so k=3/2
:
the slope of y=3/5x is 3/5